How long will it take $750 to quadruple in an account that pays 4.5% interest compounded continuously?
Answer by lwsshak3(11628)

How long will it take $750 to quadruple in an account that pays 4.5% interest compounded continuously?
**
Formula for continuous compounding:A= Pe^rt, P=initial investment, r=interest per period, t=number of periods, A=amount after t periods.
For given problem:
P=750
r=4.5% per year
t=number of years
A=4P
..
A/P=e^rt
4=e^rt=e^.045t
take log of both sides
.045t*lne=ln4
log of base=lne=1
.045t=ln4
t=ln4/.045
t=30.8
ans:
It would take about 40 years to quadruple in an account that pays 4.5% interest compounded continuously?
note: this would be true for any initial investment compounded continuously at 4.5% interest rate.